OpenStudy (anonymous):

The expression 3/x-1 - 6 will equal 0 when x equals which of the following?

OpenStudy (atlas):

There are several ways in which you can solve this problem

OpenStudy (anonymous):

The equation is: \[\frac{ 3 }{ x-1 } - 6\]

OpenStudy (gorv):

3/(x-1) -6=0 3/(x-1)=6 1/(x-1)=2 1/2=x-1 x=1+1/2 x=3/2=1.5

OpenStudy (gorv):

@SlightlyConfused got it????

OpenStudy (atlas):

One way is to know that when two sides are equal - and if you operate both sides in the same way they still remain equal.

OpenStudy (atlas):

So we have: \[\frac{ 3 }{ x-1 } -6 =0\] I shall add 6 on both sides and they still remain equal

OpenStudy (atlas):

\[\frac{3 }{ x-1 } =6\]

OpenStudy (atlas):

Now i multiply both sides by (x-1)

OpenStudy (atlas):

3 = 6(x-1) I guess you can solve for x after this

OpenStudy (anonymous):

ohhhhh okay i got it!!!!!!!!!!!!!

OpenStudy (anonymous):

9/6 ??

OpenStudy (atlas):

Yeah .........check if you can simplify 9/6 further

OpenStudy (anonymous):

3/2 !!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (atlas):

right !!!

OpenStudy (anonymous):

Thanks @atlas @gorv